OpenStudy (anonymous):

Solve 2x^2 - 8x = -7. I think by completing the square

OpenStudy (anonymous):

@Vijeya3 any thoughts?

OpenStudy (anonymous):

oh alright thank you so much !

jhonyy9 (jhonyy9):

@Everly what is the first step ?

OpenStudy (anonymous):

you take the square root right?

jhonyy9 (jhonyy9):

no i think take the -7 on the left side and how will get than the quadratic ?

OpenStudy (anonymous):

take the -7 and do what?

jhonyy9 (jhonyy9):

just than see what you get ?

jhonyy9 (jhonyy9):

@Vijeya3 Attention please here in openstudy we dont give to asker directly right answer just we try help step by step - ok ?

OpenStudy (anonymous):

I'm still unclear of which route to take, I don't know how to get to the trinomial either

jhonyy9 (jhonyy9):

@mathmale your opinion please about this ?

OpenStudy (anonymous):

I understand taking the square root, completing the square, quad formula etc

OpenStudy (anonymous):

im going to take the quadratic formula i think, and yes ax^2+bx+c=0, right?

OpenStudy (anonymous):

i think so but doesn't the x^2 coefficient have to equal to 1?

OpenStudy (anonymous):

i tried to divide 2 but that left me with a fraction i didnt know how to solve

OpenStudy (anonymous):

are we starting from before or after we add 7 to both sides?

OpenStudy (anonymous):

yeah to make it equal to 0

OpenStudy (anonymous):

a quadratic formula? 2x^2 - 8x +7 = 0

OpenStudy (anonymous):

yes i did it a while ago but forgot it, (x+4) (x-4) ??

OpenStudy (anonymous):

so can you explain the method then?

OpenStudy (anonymous):

x≈2.70710678,1.29289322

OpenStudy (anonymous):

I understand all of that trust me I do and I've done other expressions but its just this one I'm stuck on

OpenStudy (anonymous):

@RhondaSommer

RhondaSommer (rhondasommer):

\[2x^2-8x + 7 = 0\] this is what you know right?

RhondaSommer (rhondasommer):

@Vijeya3 i was personally asked to help. please do not interfere while I teach.

OpenStudy (anonymous):

Yes, but I don't know where to go after it. I can't use the sum-product, it doesn't make sense

OpenStudy (anonymous):

medal?

RhondaSommer (rhondasommer):

okay Divide each side by '2'.

RhondaSommer (rhondasommer):

sorry each coefficient

RhondaSommer (rhondasommer):

because you want the coefficient of x^2 to be 1

OpenStudy (anonymous):

i did that but i ended up with the fraction 7/2 (3.5) x^2+4+7/2 (3.5) = 0

RhondaSommer (rhondasommer):

yup :) you need that to continue

RhondaSommer (rhondasommer):

\[ x^2 - 4x + 3.5= 0\]

RhondaSommer (rhondasommer):

now; you need to subtract 3.5 from both sides :)

OpenStudy (anonymous):

Really? I was trying to complete the square... okay so x^2 +4 = 3.5 what do I do next?

RhondaSommer (rhondasommer):

\[x^2 + 4 = -3.5\] not just 3.5 because you subtracted

OpenStudy (anonymous):

Oh sorry, I wrote that but didn't type, do I factor?

RhondaSommer (rhondasommer):

So now your x term is -4x. Take half its coefficient (-2). Square it (4) and add it to both sides.

RhondaSommer (rhondasommer):

\[x^2 + 4 + 4 = -3.5 +4\]

OpenStudy (anonymous):

How did you know to do that? What method is that?

RhondaSommer (rhondasommer):

completing the square. \[x^2 + 4x + 4 = -3.5 + 4\]

OpenStudy (anonymous):

That's the part I'm confused at, because in order to complete square the coefficient of x should be equal to 2a

OpenStudy (anonymous):

Which I know see is 2(2) = 4

OpenStudy (anonymous):

now*

RhondaSommer (rhondasommer):

and our x coefficient is 2 so does that make sense now?

OpenStudy (anonymous):

Yup! So add 4 to both sides and get x^2 +8 = 7.5, correct?

RhondaSommer (rhondasommer):

-3.5 + 4 = 0.5. i also miss wrote it at first \[x^2+4x+4=-3.5+4\]

OpenStudy (anonymous):

This is making so much sense now thank you, okay so , I subtract for now from both sides?

OpenStudy (anonymous):

four*

RhondaSommer (rhondasommer):

nope simplify. ^_^ \[x^2+4x+4=0.5\]\[x^2+4x+4\] is a perfect square...so simplify it ^_^ then tell me what it is :)

OpenStudy (anonymous):

Yup! I just caught on, x^2 + 4x + 4 (x+2)^2 I think, not too sure

RhondaSommer (rhondasommer):

whoops! i keep writing it wrong. my mistake. \[x^2-4x+4\] simplify that

RhondaSommer (rhondasommer):

i made that mistake waaay earlier. it should have been -4x the entire time. luckily that doesnt affect what we have done so far.

RhondaSommer (rhondasommer):

instead; it will be (x-2)(x-2) = the square root of ±0.5

OpenStudy (anonymous):

but would that four we added now have to be subtracted x^2 - 4x -4 ?

RhondaSommer (rhondasommer):

no because that is a perfect square ^_^

RhondaSommer (rhondasommer):

(x-2) = -0.70710678118654752440084436210485 and (x-2) = 0.70710678118654752440084436210485

RhondaSommer (rhondasommer):

simplify those :)

OpenStudy (anonymous):

So 2(a) is always 2(a) no matter if x is a negative ?

RhondaSommer (rhondasommer):

yup :)

RhondaSommer (rhondasommer):

1.2928932188134524755991556378952 2.7071067811865475244008443621048 when simplified :) does that make sense?

OpenStudy (anonymous):

Sorry my computer froze, I'm not sure if I understand how to simplify those. And my options are way different too!

RhondaSommer (rhondasommer):

thats strange. what are your options

RhondaSommer (rhondasommer):

simplify it :)

RhondaSommer (rhondasommer):

1.4142135623730950488016887242097 is square root of 2

RhondaSommer (rhondasommer):

and it has to be positive 2 not negative because of (x-2) add two and thats part of your answer

RhondaSommer (rhondasommer):

so i belive it is d :) √2 /2 is the same as √0.5

RhondaSommer (rhondasommer):

:) make sense?

OpenStudy (anonymous):

I kind of understand that but what do you mean by (x-2) add 2 is a part of my answer?

RhondaSommer (rhondasommer):

no. x-2 = 0 you have to add 2 two both sides to get x alone

RhondaSommer (rhondasommer):

its just to get x alone. your answer is simply d

RhondaSommer (rhondasommer):

::)

OpenStudy (anonymous):

I kind of understand that but what do you mean by (x-2) add 2 is a part of my answer?

OpenStudy (anonymous):

Oh!!!! Wow I finally get it! I've been on this question since 9 AM and it's now almost 11:30AM. I couldn't figure it out without you @RhondaSommer thank you so so so much :)

RhondaSommer (rhondasommer):

no problemo ^_^